please check if im right!!

(2x^3y^-2)^-2 = y^4/4x^6
-5x^2y^4z^6/45x^7y^-8z^4 = -y^12z^2/9x^5
6.2 * 10^4/3.1*10^-2 =2.6*10^6 = 2000000
(3.25*10^-6)*(2.4*10^2) = 7.8*10^4
M= cubed root 1-p, when p=28
is M = to -3?
Thanks in advance!

Answer by oscargut(2103)

(2x^3y^-2)^-2 = y^4/4x^6 (you are right)
-5x^2y^4z^6/45x^7y^-8z^4 = -y^12z^2/9x^5 (you are right)
6.2 * 10^4/3.1*10^-2 =2.6*10^6 = 2000000 (is 2*10^6)
(3.25*10^-6)*(2.4*10^2) = 7.8*10^4 (is 7.8*10^-4)
M= cubed root 1-p, when p=28
is M = to -3? (yes, M=-3)